Find the regression lines (both) from the following data:
X: 1 2 3 4 5 6
Y: 3 5 6 8 9 11
Also find y for x = 2.5 and x for y=7
1
Expert's answer
2020-08-20T18:12:58-0400

From the given data of x & y


CASE 1 :


Let us consider that y is the dependent variable and x is an independent variable then the formula for the regression line is given by,


y = a + bx


where,

a = y intercept

b = slope of the line


The formula for a and b is given by,


"b = \frac{n\sum(xy)- \sum x\sum y}{n\sum(x)^2-(\sum x)^2}

\ \ \ \ \ \ \ and\ \ \ \ \ 

a = \frac{\sum y}{n} - b\frac{\sum x}{n}"


where n = sample size = 6




from the above table we get the values of a and b as under


"b = \frac{n\sum(xy)- \sum x\sum y}{n\sum(x^2)-(\sum x)^2} = \frac{6*174 - 21*42}{6*91 - 21^2} = 1.5428"


"a = \frac{\sum y}{n} - b\frac{\sum x}{n} = \frac{42}{6} - 1.542857*\frac{21}{6} = 1.6"


Hence the equation of the regression line becomes y = 1.6 + 1.54x


So the value of y, when x = 2.5 can be found using above equation

y = 1.6 + 1.54*2.5 = 5.457


(X1,Y1) = (2.5 , 5.457)


CASE 2 :


Let us consider that x is the dependent variable and y is an independent variable then the formula for the regression line is given by,


x = a + by


where,

a = x intercept

b = slope of the line


The formula for a and b is given by,


"b = \frac{n\sum(xy)- \sum x\sum y}{n\sum(y)^2-(\sum y)^2}

\ \ \ \ \ \ \ and\ \ \ \ \ 

a = \frac{\sum x}{n} - b\frac{\sum y}{n}"


where n = sample size = 6




from the above table we get the values of a and b as under


"b = \frac{n\sum(xy)- \sum x\sum y}{n\sum(y)^2-(\sum y)^2} = \frac{6*174 - 21*42}{6*336-42^2} = \frac{9}{14} = 0.64285"


"a = \frac{\sum x}{n} - b\frac{\sum y}{n} = \frac{21}{6} - 0.64285*\frac{42}{6} = -1"


Hence the equation of the regression line becomes x = -1 + 0.64y


So the value of x, when y = 7 can be found using above equation

x = -1 + 0.64*7 = 3.48


(X2 , Y2) = (3.48 , 7)